I need your help I will appreciate it.
The Great Pyramid of Giza in Cairo,Egypt is a square pyramid. Each edge of the base is 751 feet and the slant height of each face is 610 feet. Find the approximate surface area of the pyramid.

Found 2 solutions by ilana, checkley71:
Answer by ilana(307)

OK, a square pyramid has 5 faces- 4 of the same triangle and 1 base. So, the surface area is the area of those 5 faces. The base has an area of 751*751=564001ft^2 and each triangle has an area of 1/2*751*610=229055ft^2.
So the surface area of the pyramid is 564001 + 4*229055 = 1480221 ft^2.

Answer by checkley71(8403)

THE SURFACE OF THIS PYRAMID IS THE SUM OF THE AREA OF ALL 4 SIDES.
EACH SIDE IS A TRIANGLE WITH A HEIGHT OF 610 & A BASE OF 751.
THE AREA OF EACH TRIANGLE IS (BASE*HEIGHT)/2.
THE NUMBER OF SIDES=4
THE SLANT HEIGHT=610
1/2 OF THE BASE MEASUREMENT=(751/2)
4*(610*751/2)=916,220 FT^2 IS THE TOTAL SURFACE AREA OF THE 4 SLANT SIDES.